# Six-Point Method for Multi-Camera Systems with Reduced Solution Space

### v1

@@ 3.4 Polynomial System Solving
Based on the polynomial equation system Eqs. (41) and (42), we propose a minimal generic solver for the generalized camera and two minimal solvers for standard configurations of two-camera rigs. The Gröbner basis technique can be used to find algebraic solutions to the polynomial equation system [28, 39]. Firstly, we construct a random instance of the original equation system in a finite prime field $\mathbb{Z}_{p}$ [36] or a rational field. This is beneficial for maintaining numerical stability and eliminating large number arithmetic during the calculation of Gröbner basis. Secondly, Macaulay 2 [11] is used to calculate Gröbner basis. Finally, we use an automatic Gröbner basis solver [28] to find the solver of the polynomial equation system. It should be noted that the polynomial equations $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ can be extended to deal with relative pose estimation for other configurations of multi-camera systems, such as with partially uncalibrated cameras and known rotation angles.
@@ 3.4.1 Minimal Solver for Generalized Cameras
We propose a generic solver for the generalized camera given six PCs. As shown in Fig. 2, there are two views of a generalized camera, and there are 6 PCs across two views. We can define 12 virtual perspective cameras by the following method. The origins $o_{1},\cdots,o_{6}$ and $o^{\prime}_{1},\cdots,o^{\prime}_{6}$ of the virtual cameras are the positions of PCs. The orientations of these virtual cameras are consistent with the generalized camera’s reference. The 6 PCs can be equivalently captured by a virtual 12-camera rig. Moreover, we can calculate the extrinsic parameters of these 12 virtual perspective cameras and the image coordinates of PCs in these virtual cameras. As a result, we can construct a minimal generic solver to recover the relative pose of the generalized camera.
Figure 2: Relative pose estimation for generalized cameras. Specifically, two-view geometry for 6 PCs and a generalized camera can be described by a 12-camera rig. Note that points $o_{i}$ and $o^{\prime}_{i}$ do not necessarily correspond to the same physical point of the generalized camera.
Tab. 1 shows the statistics of the proposed minimal solvers for generalized cameras. #sym represents the number of symmetries, #sol represents the number of solutions, and $1$-dim represents one dimensional extraneous roots. The Cayley parameterization and quaternion parameterization solvers are named as 6pt+cayl+generic and 6pt+quat+generic, respectively. The observations are summarized as follows: (1) $\mathcal{E}_{2}$ is an empty set, and $\mathcal{E}_{1}$ is sufficient to solve the relative pose. (2) Due to one-fold symmetry in quaternion, Cayley parameterization results in fewer solutions than quaternion parameterization. The number of complex solutions obtained by the 6pt+cayl+generic solver and 6pt+quat+generic solver is 64 and 128, respectively. (3) Elimination templates of the 6pt+cayl+generic solver and the 6pt+quat+generic solver are $99\times 163$ and $342\times 406$, respectively. Given that the solvers using Cayley parameterization yields smaller eliminate templates compared to the solvers using quaternion parameterization, we adopt the former as our default choice in this paper.
@@ 3.4.2 Minimal Solvers for Two-camera Rigs
We propose two minimal solvers for two practical configurations of two-camera rigs in Fig. 3. These solvers comprise an inter-camera solver and an intra-camera solver, and both configurations offer two ray bundle constraints within $\mathcal{E}_{2}$. The inter-camera solver utilizes inter-camera PCs that are observable to different cameras across two views. This solver is appropriate for multi-camera systems with significant overlapping views. The intra-camera solver utilizes intra-camera PCs that are observable to the same camera across two views. It is appropriate for multi-camera systems with non-overlapping or small overlapping views.
For inter-camera case, $\mathcal{E}_{1}$ is sufficient to solve the relative pose of two-camera rigs. The number of solutions can be reduced by employing both $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$. For intra-camera case, one-dimensional families of extraneous roots exist when only $\mathcal{E}_{1}$ is employed. Combining $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ can solve the relative pose in the intra-camera case. The solvers for these two cases using Cayley parametrization are named as 6pt+cayl+inter and 6pt+cayl+intra, respectively. The solvers for these two cases using quaternion parameterization are named as 6pt+quat+inter and 6pt+quat+intra, respectively.
(a) Inter-camera case.
(b) Intra-camera case.
Tab. 1 also shows the statistics of the proposed minimal solvers for two-camera rigs. The observations are summarized as follows: (1) When $\mathcal{E}_{1}$ is employed, the inter-camera solver has up to $56$ complex solutions, and there are one-dimensional families of extraneous roots in the intra-camera case. (2) By employing both $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$, the number of complex solutions obtained by both the inter-camera and intra-camera solvers is $48$. (3) Compared to the solvers using quaternion parameterization, the solvers using Cayley parameterization yield smaller eliminate templates. We adopt Cayley parameterization as our default choice. (4) When only $\mathcal{E}_{1}$ is used in the 6pt+quat+inter configuration, it is necessary to explicitly consider the inequality $q_{w}\neq 0$. Otherwise, one-dimensional extraneous roots exist. We account for this inequality utilizing the saturation method [29], which yields $112$ solutions with one-fold symmetry. (5) For the 6pt+cayl+inter configuration, compared to the solvers derived from both $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$, the solvers solely derived from $\mathcal{E}_{1}$ yield smaller eliminate templates and exhibit better numerical stability. This phenomenon indicates that the number of basis might affect the numerical stability of the solvers, which has been previously observed in the literatures [5, 13].
@@ 4 Experiments
Runtime 43.3 102.0 3275.4 2447.4 1677.8 1449.4 1463.8
(a) Rotation error $\varepsilon_{\mathbf{R}}$.
(b) Translation error $\varepsilon_{\mathbf{t}}$.
(c) Translation direction error $\varepsilon_{\mathbf{t},\text{dir}}$.
Figure 4: Probability density functions of the relative pose estimation error on noise-free observations. We designed a simulated two-camera rig comprising two perspective cameras. The orientations of two perspective cameras are roughly forward-facing with random perturbation. The horizontal axis corresponds to the $\log_{10}$ estimated errors, and the vertical axis is the empirical probability density.
@@
Figure 5: Relative pose estimation error with varying image noise for a generalized camera. We design a simulated multi-camera system comprising omnidirectional cameras. The extrinsic parameters, including orientation and position, are totally random. The three rows correspond to the generic, inter-camera, and intra-camera cases, respectively. The 6pt-Our solver in the three rows represent 6pt-Our-generic, 6pt-Our-inter, and 6pt-Our-intra, respectively.
This section presents a series of experiments performed on synthetic and real-world datasets to evaluate the performance of the proposed solvers. All the proposed solvers employ Cayley parameterization for their implementations. The solver presented in Sec. 3.4.1 is named as the 6pt-Our-generic method. The solvers proposed in Sec. 3.4.2, which aim to handle inter-camera and intra-camera cases, are named as the 6pt-Our-inter and 6pt-Our-intra solvers, respectively. To further differentiate between different solvers for 6pt-Our-inter, we denote the solvers resulting from ${\mathcal{E}}_{1}$ and ${\mathcal{E}}_{1}\cup{\mathcal{E}}_{2}$ as 6pt-Our-inter56 and 6pt-Our-inter48, respectively. Refer to [24, 13], the proposed solvers are evaluated against state-of-the-art solvers using PCs, including 17pt-Li [35], 8pt-Kneip [24], and 6pt-Stewénius [46]. The paper does not evaluate solvers that exploit the additional affine parameters besides PCs [15, 13] or utilize a prior for relative pose estimation [49, 50]. The proposed minimal solvers are implemented in C++ [^1] Source codes are provided in the supplementary material and will be released.. The source codes for the comparison methods are adopted from the OpenGV library [23].
In the experiments presented in Sec. 4.2 and Sec. 4.3, each solver is independently integrated into RANSAC [9] to reject outliers and obtain the estimated relative pose with the highest number of inliers. The angular re-projection error [23, 33] introduced by PCs is used to classify inliers. We follow the default parameters of OpenGV to set the inlier threshold angle as $0.1^{\circ}$ [23]. During RANSAC iterations, the outlier ratio is from the current best model. The stopping criterion of RANSAC iterations is that at least one outlier-free set is sampled with a probability $0.99$, or the maximum number $20,000$ of iterations is reached. We compute the rotation error as the angular difference between the ground truth rotation $\mathbf{R}_{\text{gt}}$ and the estimated rotation ${\mathbf{R}}$: ${\varepsilon_{\mathbf{R}}}=\arccos((\operatorname*{trace}({\mathbf{R}_{\text{gt}}}{{\mathbf{R}^{T}}})-1)/2)$. To calculate the translation error, we employ the definition introduced in [43, 32]: ${\varepsilon_{\mathbf{t}}}=2\left\|{{\mathbf{t}_{\text{gt}}}}-{\mathbf{t}}\right\|/(\left\|{\mathbf{t}_{\text{gt}}}\right\|+\left\|{{\mathbf{t}}}\right\|)$, where $\mathbf{t}_{\text{gt}}$ and ${\mathbf{t}}$ are the ground truth translation and the estimated translation, respectively. We also compute the translation direction error: $\varepsilon_{\mathbf{t},\text{dir}}=\arccos(\mathbf{t}_{\text{gt}}^{T}\mathbf{t}/(\|\mathbf{t}_{\text{gt}}\|\cdot\|\mathbf{t}\|))$.
The principles of minimal solver choice are listed below. First, when 6pt-Our-inter or 6pt-Our-intra are applicable, we apply them with higher priority than 6pt-Our-generic. The reason is that the solvers of specific configurations usually have better performance than the generic solver. Second, we recommend 6pt-Our-intra as the default solver for real-world image sequences. The relative pose is usually estimated for consecutive image pairs captured in a small time interval. Thus, intra-camera PCs can be built up with a higher probability than inter-camera PCs.
4.1 Efficiency and Numerical Stability
The runtimes of the solvers for multi-camera systems are assessed using an Intel(R) Core(TM) i7-7800X 3.50 GHz processor. The average runtime of the solvers over $10,000$ runs is shown in Tab. 2. As a linear solver, 17pt-Li is the most efficient one. All the proposed solvers have runtimes of $1.5\sim 2.5$ milliseconds, which are more efficient than the seminal 6pt-Stewénius among the minimal solvers.
To test the numerical stability, we design two simulated scenarios. First, we design a simulated two-camera rig composed of two perspective cameras. The orientations of the two perspective cameras are roughly forward-facing with random perturbation. This setting is practical for autonomous driving with two front-facing cameras. Second, we design a simulated generalized camera composed of omnidirectional cameras. The extrinsic parameters, including orientation and position, are totally random. We report the results for the solvers under the first scenario below. The results for generalized cameras under the second scenario are shown in the supplementary material.
Under the first scenario, the numerical stability of the solvers on noise-free observations is shown in Fig. 4. The procedure is repeated for $10,000$ times on both inter-camera and intra-camera cases. The resulting empirical probability density functions are represented as a function of the $\log_{10}$ estimated errors. The numerical stability of a solver is determined by many factors, such as the problem’s complexity, parameterization, equation system construction, solver generator, and implementation. Usually, more efficient solvers generated by effective methods have less round-off error and thus have better numerical stability. Among these solvers, the 17pt-Li solver demonstrates the best numerical stability due to its status as a linear solver with fewer computations. Since the 8pt-Kneip solver relies on iterative optimization, it is susceptible to becoming trapped in local minima. All the proposed solvers exhibit satisfactory numerical stability. Since the 6pt-Our-inter56 solver displays better numerical stability than the 6pt-Our-inter48 solver, we recommend 6pt-Our-inter56 as the default solver for inter-camera case in the follow-up experiments.
4.2 Experiments on Synthetic Data
Following Sec. 4.1, two simulated scenarios are designed and tested for the synthetic experiments, including scenarios for a two-camera rig and a generalized camera. The results for two-camera rigs under the first scenario are shown in the supplementary material.
@@ 4.2 Experiments on Synthetic Data
Under the second scenario, the simulated scenario for a generalized camera is described as below. For the generic case, a generalized camera comprises 12 omnidirectional cameras. For the inter-camera and intra-camera cases, a generalized camera comprises 2 omnidirectional cameras. The extrinsic parameters of each omnidirectional camera are generated randomly. Each scene point is generated randomly. In addition, the relative poses between the two views are also generated randomly. Omnidirectional cameras are selected in our settings, because usually there is no overlap for pinhole cameras with random extrinsic parameters and relative poses. We test the accuracy of pose estimation for all the proposed solvers, including 6pt-Our-generic, 6pt-Our-inter, and 6pt-Our-intra solvers.
@@ 4.3 Experiments on Real-World Data
To evaluate the performance of the proposed solvers in practical applications, three public datasets KITTI [10], nuScenes [6], and EuRoc [4] are used in the experiments. Specifically, the KITTI and nuScenes datasets are collected in autonomous driving environments, while the EuRoc dataset is collected in a micro aerial vehicle environment. These datasets contain challenging image pairs, such as highly dynamic scenes and significant motion. The proposed solvers are compared with state-of-the-art solvers. The accuracy of relative pose estimation is evaluated using the rotation error ${\varepsilon_{\bf{R}}}$ and the translation direction error $\varepsilon_{\mathbf{t},\text{dir}}$ [24, 37]. Our evaluation is based on approximately $30,000$ image pairs, and we report the final estimation results by integrating the minimal solver with RANSAC. The results for nuScenes and EuRoc datasets are provided in the supplementary material.
@@
Mean time 3.157 0.648 4.723 1.849
Std. deviation 0.119 0.009 0.161 0.072
We evaluate all the solvers on KITTI dataset [10], which is collected using outdoor autonomous vehicles installed with forward-facing stereo cameras. It is treated as a general multi-camera system by ignoring overlapping fields of view for cameras. The 6pt-Our-intra solver is tested on 11 available sequences containing 23,000 image pairs. The ground truth is obtained directly from the GPS/IMU localization unit [10]. To establish PCs for consecutive views of each camera, the SIFT method [38] is used. Additionally, all the solvers are integrated into RANSAC to remove mismatches in the experiments.
Tab. 3 illustrates the rotation and translation error of the 6pt-Our-intra solver on KITTI dataset. We use median error to evaluate the performance of solvers. The proposed 6pt-Our-intra solver outperforms the comparative solvers in overall performance. To further compare computational efficiency, Tab. 4 illustrates the corresponding average runtime of RANSAC on KITTI dataset. Although 17pt-Li is faster than the proposed 6pt-Our-intra, the proposed solver demonstrates better efficiency when integrated them separately into the RANSAC framework.

### v2

@@ 3.4 Polynomial System Solving
Based on the polynomial equation system Eqs. (14) and (15), we propose a minimal generic solver for the generalized camera and two minimal solvers for standard configurations of two-camera rigs. The Gröbner basis technique can be applied to discover algebraic solutions for the polynomial equation system [28, 40]. Firstly, we construct a random instance of the original equation system in either a finite prime field $\mathbb{Z}_{p}$ [37] or a rational field. This strategy helps to maintain numerical stability and avoid arithmetic with large numbers when computing Gröbner basis. Secondly, Macaulay 2 [11] is utilized for computing Gröbner basis. Finally, we use an automatic Gröbner basis solver [28] to find the solution of the polynomial equation system. It should be noted that the polynomial equations $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ can be extended to deal with relative pose estimation for other configurations of multi-camera systems, such as with partially uncalibrated cameras and known rotation angles.
@@ 3.4.1 Minimal Solver for Generalized Cameras
We propose a generic solver for the generalized camera given six PCs. As shown in Fig. 3, there are two views of a generalized camera, and there are 6 PCs across two views. Specifically, two-view geometry for 6 PCs and a generalized camera can be described by a 12-camera rig. We can define 12 virtual perspective cameras by the following method. The origins $o_{1},\cdots,o_{6}$ and $o^{\prime}_{1},\cdots,o^{\prime}_{6}$ of the virtual cameras are the positions of PCs. The orientations of these virtual cameras are consistent with the generalized camera’s reference. The 6 PCs can be equivalently captured by a virtual 12-camera rig. Moreover, we can calculate the extrinsic parameters of these 12 virtual perspective cameras and the image coordinates of PCs in these virtual cameras. Please see the supplementary material for details. As a result, we can construct a minimal generic solver to recover the relative pose of generalized cameras.
Figure 2: Relative pose estimation for generalized cameras. Note that points $o_{i}$ and $o^{\prime}_{i}$ do not necessarily correspond to the same physical point of the generalized camera.
Tab. 1 shows the statistics of the proposed minimal solvers for generalized cameras. Here, #sym indicates the number of symmetries, #sol indicates the number of solutions, and $1$-dim indicates one-dimensional extraneous roots. The Cayley parameterization and quaternion parameterization solvers are named as 6pt+cayl+generic and 6pt+quat+generic, respectively. The observations are summarized as follows: (1) $\mathcal{E}_{2}$ is an empty set, and $\mathcal{E}_{1}$ is sufficient to solve the relative pose. (2) Due to one-fold symmetry in quaternion, Cayley parameterization results in fewer solutions than quaternion parameterization. The number of complex solutions yielded by the 6pt+cayl+generic solver and 6pt+quat+generic solver is 64 and 128, respectively. (3) Elimination templates of the 6pt+cayl+generic solver and the 6pt+quat+generic solver are $99\times 163$ and $342\times 406$, respectively. Given that the solvers using Cayley parameterization yields smaller eliminate templates compared to the solvers using quaternion parameterization, we adopt the former as our default choice in this paper.
@@ 3.4.2 Minimal Solvers for Two-camera Rigs
Two minimal solvers are proposed for two practical configurations of two-camera rigs in Fig. 3. These solvers comprise an inter-camera solver and an intra-camera solver, and both configurations offer two ray bundle constraints within $\mathcal{E}_{2}$. The inter-camera solver utilizes inter-camera PCs that are observable to different cameras across two views. This solver is appropriate for multi-camera systems characterized by significant overlap between views. Conversely, the intra-camera solver utilizes intra-camera PCs that are observable to the same camera across two views. This solver is appropriate for multi-camera systems characterized by small or no overlap between views.
For inter-camera case, $\mathcal{E}_{1}$ is enough to compute the relative pose of two-camera rigs. The number of solutions can be reduced by employing both $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$. For intra-camera case, one-dimensional families of extraneous roots exist when only $\mathcal{E}_{1}$ is employed. Combining $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ can solve the relative pose in the intra-camera case. The solvers for these two cases using Cayley parametrization are named as 6pt+cayl+inter and 6pt+cayl+intra, respectively. The solvers for these two cases using quaternion parameterization are named as 6pt+quat+inter and 6pt+quat+intra, respectively.
Tab. 1 also shows the statistics of our minimal solvers for two-camera rigs. The observations are summarized as follows: (1) When $\mathcal{E}_{1}$ is employed, the inter-camera solver has up to $56$ complex solutions, and there exist one-dimensional families of extraneous roots in the intra-camera case. (2) By employing both $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$, both the inter-camera and intra-camera solvers yield a total of 48 complex solutions. (3) Compared to the solvers using quaternion parameterization, the solvers using Cayley parameterization yield smaller eliminate templates. We adopt Cayley parameterization as our default choice. (4) When only $\mathcal{E}_{1}$ is used in the 6pt+quat+inter configuration, it is necessary to explicitly consider the inequality $q_{w}\neq 0$. Otherwise, one-dimensional extraneous roots exist. We account for this inequality utilizing the saturation method [29], which yields $112$ solutions exhibiting one-fold symmetry. (5) For the 6pt+cayl+inter configuration, compared to the solvers derived from both $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$, the solvers solely derived from $\mathcal{E}_{1}$ yield smaller eliminate templates and exhibit better numerical stability. This phenomenon indicates that the number of bases might impact the numerical stability of the solvers, which has been previously observed in the literature [5, 14].
@@ 4 Experiments
This section presents a series of experiments performed on synthetic and real-world datasets to assess the performance of our solvers. All the proposed solvers employ Cayley parameterization for their implementations. The solver for the generalized camera is named as the 6pt-Our-generic method. The solvers aim to handle inter-camera and intra-camera cases, are named as the 6pt-Our-inter and 6pt-Our-intra solvers, respectively. To further differentiate between different solvers for 6pt-Our-inter, we denote the solvers resulting from ${\mathcal{E}}_{1}$ and ${\mathcal{E}}_{1}\cup{\mathcal{E}}_{2}$ as 6pt-Our-inter56 and 6pt-Our-inter48, respectively. Following [24, 14], our solvers are evaluated against state-of-the-art solvers using PCs, including 17pt-Li [36], 8pt-Kneip [24], and 6pt-Stewénius [48]. The paper does not evaluate solvers that exploit the additional affine parameters besides PCs [15, 14] or utilize a prior for relative pose estimation [50, 51]. The proposed minimal solvers are implemented in C++. The source codes for 17pt-Li [36] and 8pt-Kneip [24] are adopted from the OpenGV library [23]. The source codes for 6pt-Stewénius [48] is adopted from the PoseLib library [30].
In the experiments presented in Sec. 4.1 and Sec. 4.2, each solver is independently integrated into RANSAC [9] to reject outliers and obtain the estimated relative pose with the highest number of inliers. To ensure the fairness of the experiment, all the different solvers are compared over the same data and within the same RANSAC implementation. In addition, we do not apply a non-minimal solver or perform optimization for the estimated relative pose with all inliers. The angular re-projection error [23, 34] induced by PCs is used to classify inliers. We follow the default parameters of OpenGV to set the inlier threshold angle as $0.1^{\circ}$ [23]. During RANSAC iterations, the outlier ratio is from the current best model. The stopping criterion of RANSAC iterations is that at least one outlier-free set is sampled with a probability $0.99$, or the maximum number $20,000$ of iterations is reached. We compute the rotation error as the angular difference between the ground truth rotation $\mathbf{R}_{\text{gt}}$ and the estimated rotation ${\mathbf{R}}$: ${\varepsilon_{\mathbf{R}}}=\arccos((\mathrm{tr}({\mathbf{R}_{\text{gt}}}{{\mathbf{R}^{T}}})-1)/2)$. To calculate the translation error, we employ the definition introduced in [44, 33]: ${\varepsilon_{\mathbf{t}}}=2\left\|{{\mathbf{t}_{\text{gt}}}}-{\mathbf{t}}\right\|/(\left\|{\mathbf{t}_{\text{gt}}}\right\|+\left\|{{\mathbf{t}}}\right\|)$, where $\mathbf{t}_{\text{gt}}$ and ${\mathbf{t}}$ are the ground truth translation and the estimated translation, respectively. We also compute the translation direction error: $\varepsilon_{\mathbf{t},\text{dir}}=\arccos(\mathbf{t}_{\text{gt}}^{T}\mathbf{t}/(\|\mathbf{t}_{\text{gt}}\|\cdot\|\mathbf{t}\|))$.
The principles of minimal solver choice are listed below. First, when 6pt-Our-inter or 6pt-Our-intra are applicable, we apply them with higher priority than 6pt-Our-generic. The reason is that the solvers of specific configurations usually have better performance than the generic solver. Second, we recommend 6pt-Our-intra as the default solver for real-world image sequences. The relative pose is usually estimated for consecutive image pairs captured in a small time interval. Thus, intra-camera PCs can be built up with a higher probability than inter-camera PCs. The results for efficiency and numerical stability are shown in the supplementary material.
4.1 Experiments on Synthetic Data
Two simulated scenarios are designed and tested for the synthetic experiments. First, we design a simulated two-camera rig composed of two perspective cameras. The orientations of the two perspective cameras are roughly forward-facing with random perturbation. This setting is practical for autonomous driving with two front-facing cameras. Second, we design a simulated generalized camera composed of 12 omnidirectional cameras. The extrinsic parameters, including orientation and position, are totally random. The results for two-camera rigs under the first scenario are shown in the supplementary material.
@@ 4.1 Experiments on Synthetic Data
Figure 4: Relative pose estimation error with varying image noise for a generalized camera. We design a simulated multi-camera system comprising 12 omnidirectional cameras. The extrinsic parameters, including orientation and position, are totally random. The three rows correspond to the generic, inter-camera, and intra-camera cases, respectively. The 6pt-Our solver in the three rows represent 6pt-Our-generic, 6pt-Our-inter, and 6pt-Our-intra, respectively.
@@
Under the second scenario, the simulated scenario for a generalized camera is described as below. For the generic case, a generalized camera comprises 12 omnidirectional cameras. For the inter-camera and intra-camera cases, we only use 2 omnidirectional cameras of the generalized camera. The extrinsic parameters of each omnidirectional camera are generated randomly. Each scene point is generated randomly. In addition, the relative poses between two views are also generated randomly. Omnidirectional cameras are selected in our settings, because usually there is no overlap for pinhole cameras with random extrinsic parameters and relative poses. We test the accuracy of pose estimation for all the proposed solvers, including 6pt-Our-generic, 6pt-Our-inter, and 6pt-Our-intra solvers.
@@ 4.2 Experiments on Real-World Data
To assess the performance of our solvers in practical applications, three datasets KITTI [10], nuScenes [6], and EuRoc [4] are used in the experiments. Specifically, the KITTI and nuScenes datasets are collected in an autonomous driving environments, while the EuRoc dataset is collected in a micro aerial vehicle environment. These datasets contain challenging image pairs with highly dynamic scenes, such as significant motion, moving objects, and varying illumination. The proposed solvers are compared against state-of-the-art solvers. The accuracy of relative pose estimation is evaluated using the rotation error ${\varepsilon_{\bf{R}}}$ and the translation direction error $\varepsilon_{\mathbf{t},\text{dir}}$ [24, 38]. Our evaluation is based on approximately $30,000$ image pairs, and we report the final estimation results by integrating the minimal solver with RANSAC. The relative pose estimation results for the nuScenes dataset are provided in the supplementary material.
@@
Mean time 3.157 0.648 4.161 1.546
Std. deviation 0.119 0.009 0.145 0.069
4.2.1 Experiments on KITTI Dataset
We evaluate the proposed solvers on KITTI dataset [10], which is collected using outdoor autonomous vehicles installed with forward-facing stereo cameras. It is treated as a general multi-camera system, disregarding overlapping overlapping fields of view for cameras. The 6pt-Our-intra solver is tested on 11 available sequences containing 23,000 image pairs. The ground truth is obtained directly from the GPS/IMU localization unit [10]. To establish PCs for consecutive views of each camera, the SIFT method [39] is used. Additionally, all the solvers have been integrated into RANSAC to remove mismatches in the experiments.
Tab. 2 illustrates the rotation and translation error of the 6pt-Our-intra solver on the KITTI dataset. We use median error to evaluate the performance of solvers. The proposed 6pt-Our-intra solver outperforms the comparative solvers in overall performance. Since 8pt-Kneip uses an identity matrix to initialize the rotation, it has acceptable results for the forward motion in autonomous driving environment. To further compare computational efficiency, Tab. 3 illustrates the corresponding average runtime of RANSAC on the KITTI dataset. Although the runtime of 17pt-Li is lower than the proposed 6pt-Our-intra, the proposed solver demonstrates better efficiency when each is integrated separately into the RANSAC framework.
4.2.2 Experiments on EuRoC Dataset
The proposed solvers are further validated within the context of an unmanned aerial vehicle (UAV) environment, leveraging the EuRoC MAV dataset [4] for the evaluation of 6DOF relative pose estimation. This dataset records data using a stereo camera mounted on a micro aerial vehicle, with sequences labeled from MH01 to MH05 collected within a large industrial machine hall. The ground truth for the relative pose is established through a nonlinear least-squares batch solution utilizing Leica position and IMU measurements. Estimating relative pose estimation in these sequences is challenging due to the unstructured and cluttered nature of the industrial environment. Consecutive image pairs with small movement are selectively thinned out, retaining only one out of every four consecutive images for relative pose estimation. Moreover, the image pairs exhibiting insufficient motion are excluded from the experiment. The proposed solvers are evaluated against state-of-the-art solvers, including 17pt-Li [36], 8pt-Kneip [24], and 6pt-Stewénius [48]. PCs in the image pair are established using the SIFT detector [39]. All solvers are integrated into the RANSAC framework to eliminate mismatches, ensuring a more robust estimation process.
Table 4: Rotation and translation error on EuRoC dataset (unit: degree).
Seq. 17pt-Li [36] 8pt-Kneip [24] 6pt-Stew. [48]  6pt-Our-intra
MH01 0.136   3.055 0.156   3.214 0.186   4.085 0.130  2.961
MH02 0.129   2.806 0.132   2.796 0.180   3.828 0.127  2.579
MH03 0.199   2.422 0.187   2.517 0.222   3.576 0.181  2.376
MH04 0.195   3.159 0.178   3.237 0.213   5.371 0.193  3.105
MH05 0.186   3.124 0.163   2.940 0.235   4.094 0.158  2.892
Tab. 4 illustrates the rotation and translation error of the 6pt-Our-intra solver on EuRoC dataset. We use median error to assess the performance of different solvers. The experiment results demonstrate that the 6pt-Our-intra solver surpasses comparative solvers, including 17pt-Li, 8pt-Kneip, and 6pt-Stewénius. This experiment confirms the suitability of the proposed 6pt-Our-intra solver for achieving accurate 6DOF relative pose estimation in the context of unmanned aerial vehicle environments.